OpenStudy (anonymous):

y = sin(4pix) , then y'=

OpenStudy (anonymous):

\[4 \pi \cos{4 \pi x} \]

OpenStudy (anonymous):

4pi cos (4pix)

OpenStudy (anonymous):

y'= 4picos(4pix)

OpenStudy (anonymous):

that's what I got ! sweet! thanks!